OpenStudy (anonymous):

a proud fisherman is photographed next to his catch, which hanging from a spring scale supported by the ceiling of an elevator. the elevator is moving straight downwards, and slowing down at the rate of 2.45m/s^2. the spring scale reads 60N. find the weight of the fish

OpenStudy (kropot72):

g = 9.81m/s^2 The resultant acceleration of the fish = 9.81 - 2.45 = 7.36m/s^2 Let the mass of the fish = m 60 = m * 7.36 m = 60/7.36kg Now you have the mass of the fish, can you find the weight?

OpenStudy (kropot72):

Weight = mass * g = (60/7.36) * 9.81 = ?

OpenStudy (kropot72):

@dmbonu Are you there?

OpenStudy (anonymous):

the weight of the fish is 79, isnt that a high?

OpenStudy (kropot72):

The weight of the fish is 79.97 Newtons. Obviously it is a big fish. The question says the fisherman is pround of it!

OpenStudy (anonymous):

thank you so much for your help i just figure it out. 60N is the upward force (the normal force), so we are meant to be looking for the downwards force. normal force= m(a+g) so 60N=m(2.45m/s^2+9.8m/s^2) m=4.89 weight=m*g weight =48N

OpenStudy (kropot72):

Sorry, my bad :( You are correct. The resultant acceleration of the fish is (2.45 + 9.81)m/s^2. Well done!

OpenStudy (anonymous):

thank you

OpenStudy (kropot72):

You're welcome :)